# Nonlinear Superconducting Magnetoelectric Effect

### v4

@@ Abstract
A supercurrent flow can induce a nonvanishing spin magnetization in noncentrosymmetric superconductors with spin-orbit interaction. Often known as the non-dissipative magnetoelectric effect, these are most commonly found at linear order in supercurrent flow. Here, we argue that a nonlinear superconducting magnetoelectric effect (NSM) can naturally manifest in altermagnet/superconductor (ALM/SC) heterostructures: NSM manifests as a spin polarization generated as a second-order response to a driving supercurrent. Strikingly, we find NSM is the leading order magnetization response in ALM/SC heterostructures and survives even in the presence of centrosymmetry; $C_{4}\mathcal{T}$ symmetry in altermagnets zeroes both the equilibrium magnetization as well as out-of-plane linear magnetoelectric response. This renders NSM a powerful electric and non-dissipative means of controlling magnetization in ALM/SC heterostructures, a promising platform for superconducting spintronics.
Introduction.—Unlike the normal metallic state, the intrinsically non-dissipative flow of current in superconductors enables the control of spins and magnetization without dissipation Levitov et al. (1985); Edelstein (1995, 2005)—a critical functionality in realizing low-power spintronics Linder and Robinson (2015); Eschrig (2015, 2011). Microscopically, this process occurs through the superconducting magnetoelectric effect (sometimes known as the Edelstein effect) Edelstein (1995, 2005); Fujimoto (2005); Tkachov (2017); He and Law (2020, 2021); Ikeda and Yanase (2020); Chirolli et al. (2022); He et al. (2019); Yuan (2023), where a supercurrent flowing through a noncentrosymmetric metal can induce a net magnetization of the carriers. In Rashba superconductors Bauer and Sigrist (2012); Smidman et al. (2017); Houzet and Meyer (2015), such magnetoelectric effects can be naturally understood through the spin-orbit interaction (SOI): spin readily couples to the motion of electrons (and their associated Cooper pairs) producing a net spin polarization in the presence of supercurrent flow.
@@
In this work, we unveil the nonlinear superconducting magnetoelectric effect (NSM) by examining second-order spin generation in response to the supercurrent. Specifically, we find a (second-order nonlinear) supercurrent-induced magnetization $\delta\bm{M}$ in two dimensions as
@@
where $a,b=x,y$ is the direction of the supercurrent, $c=x,y,z$, and $\bm{q}$ is the Cooper pair momentum in the presence of supercurrent flow. Note $\chi_{ab}^{c}$ vanishes in the presence of $\mathcal{T}$ symmetry. As a result, to realize the second-order nonlinear response, $\mathcal{T}$ symmetry breaking is necessary. Interestingly, Eq. (1) persists even in centrosymmetric systems in sharp contrast to the case of the linear magnetoelectric effect in noncentrosymmetric superconductors He and Law (2020). As we argue, NSM can be naturally realized in magnet/superconductor heterostructures by utilizing the proximity effect between a magnetic material and an $s$-wave superconductor Fu and Kane (2008); Qi et al. (2010); Wang et al. (2015); Tkachov (2017).
For concreteness, we propose an altermagnet/superconductor (ALM/SC) heterostructure as depicted in Fig. 1(a) as a natural candidate for NSM. Altermagnets are collinear antiferromagnets with unconventional magnetic order Ma et al. (2021); Šmejkal et al. (2022a); Lee et al. (2024); Mazin et al. (2022); Mazin (2023); Feng et al. (2022); Šmejkal et al. (2022b, 2023); Ghorashi et al. (2023); Fang et al. (2023); Krempaskỳ et al. (2024); recently, several potential altermagnetic materials have been identified by ab initio simulations Ma et al. (2021); Šmejkal et al. (2022a); Lee et al. (2024). A characteristic property of $d$-wave altermagnets is $C_{4}\mathcal{T}$ symmetry, which breaks $\mathcal{T}$ and four-fold rotational ($C_{4}$) symmetry individually but preserve the combination of them. The interplay between the altermagnets and superconductivity has been theoretically investigated Zhu et al. (2023); Beenakker and Vakhtel (2023); Zhang et al. (2024); Ouassou et al. (2023); Papaj (2023); Li and Liu (2023); Giil et al. (2024).
For ALM/SC heterostructures, we find that a nonvanishing out-of-plane spin magnetization can be generated in the second-order response to the driving supercurrent. This is in sharp contrast to the case of polar or gyrotropic superconductors Edelstein (1995); He et al. (2019); He and Law (2020). This work is organized as follows: First we present a general theory of supercurrent-induced spin generation including both the linear (first order) and nonlinear (second order) terms. Then we focus on the case in a specific ALM/SC realization. Remarkably, due to $C_{4}\mathcal{T}$ symmetry in altermagnets, we find that the second-order response $\chi_{ab}^{z}$ provides the leading order contribution to the out-of-plane spin magnetization; both equilibrium out-of-plane magnetization and linear-order out-of-plane magnetoelectric effects vanish. Additionally, due to the presence of a weak SOI in real altermagnets, the linear magnetoelectric effect also shows up in the in-plane direction. Furthermore, in the weak oscillating regime, driving an AC supercurrent can induce the second harmonic and rectified magnetization. Finally, we discuss possible material candidates that include thin films of the altermagnet RuO2 and KRu4O8 on an $s$-wave superconductor. We propose that ALM/SC devices allow for the non-dissipative electric control of magnetization, which is important for high density magnetic memories and offers tantalizing possibilities for spin transport in which Joule heating and dissipation are minimized.
@@
Superconducting magnetoelectric effect—We first examine the superconducting magnetoelectric effect within an effective Bogoliubov-de Gennes (BdG) framework. We note that our treatment is general and is agnostic to the precise microscopics of the superconducting gap: the superconducting state can be intrinsic or proximity induced. In what follows, we describe either case phenomenologically via the pairing potential $\Delta_{\bm{k}}$. The finite-$\bm{q}$ BdG Hamiltonian in the Nambu basis $(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})$ reads
@@
where $H_{\bm{k}}$ is the Bloch Hamiltonian of the normal state, and $\bm{q}$ is the momentum of the Cooper pair. As we will see below, altermagnetic $H_{\bm{k}}$ enables the NSM effect. For simplicity we have used a $\bm{q}$-independent gap function $\hat{\Delta}_{\bm{k}}$, which is valid for weak values of the applied current Chirolli et al. (2022).
Table 1: Symmetry restrictions for three components of $\bm{M}$ in a two-dimensional system. Note that $C_{4}\mathcal{T}$ corresponds to the case for altermagnets that we focus on in our work. 1(2) indicates first (second) order response. $\checkmark$ indicates allowed and $\times$ indicates forbidden.
$\mathcal{P}$ $\mathcal{T}$ $\mathcal{PT}$ $C_{4}\mathcal{T}$
$M_{x}$(2) ✓(1) $\times$ ✓(1)
$M_{y}$(2) ✓(1) $\times$ ✓(1)
$M_{z}$(2) ✓(1) $\times$ ✓(2)
The total free energy $\mathcal{F}$ of the ALM/SC heterostructure comprises two parts: the superconductor component $\mathcal{F}_{\mathrm{sc}}$ and the material (i.e. altermagnet) component $\mathcal{F}_{\mathrm{al}}$ so that $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{\mathrm{al}}$. The supercurrent can be tracked as $\bm{J}_{s}=\nu_{s}\bm{q}$ with the condensate superfluid stiffness $\nu_{s}=\partial^{2}\mathcal{F}_{\mathrm{sc}}(\bm{q})/\partial^{2}\bm{q}$. As a result, observables that scale with $q$ below are induced by the supercurrent at linear order while $q^{2}$ indicates a second-order nonlinear response.
The magnetic response of the heterostructure can be obtained in the standard fashion by analyzing the set of parametric partition function $\mathcal{Z}(\bm{h})$
@@
where $\bm{h}$ is a parameter describing an auxiliary Zeeman field with a corresponding Zeeman energy $\Sigma(\bm{h})=\mathrm{diag}(g_{s}\mu_{B}\bm{h}\cdot\hat{s},-g_{s}\mu_{B}\bm{h}\cdot\hat{s}^{*})/2$. Here $g_{s}=2$ is the Landé $g$ factor, $\mu_{B}$ being the Bohr magneton, and the Gor’kov’s Green’s functions are $G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})=\mathcal{G}^{-1}(\bm{k},\bm{q},i\omega_{n})+\Sigma(\bm{h})$ and $\mathcal{G}(\bm{k},\bm{q},i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}})^{-1}$ with the Matsubara frequency $\omega_{n}=(2n+1)\pi/\beta$. $\beta=1/k_{B}T$ with $T$ being the temperature.
@@
Here $\eta_{a}=\mathrm{diag}(s_{a},-s_{a}^{*})$ are the generalized spin Pauli matrices in the Nambu space. To extract the linear and nonlinear responses systematically, we expand the BdG Hamiltonian as
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}=0}+\frac{1}{2}q_{a}\hat{v}_{a}+\frac{1}{8}q_{a}q_{b}\hat{w}_{ab},$ (5)
where $\hat{v}_{a}=\mathrm{diag}[\mathcal{V}_{a}(\bm{k}),-\mathcal{V}_{a}^{*}(-\bm{k})]$ and $\hat{w}_{ab}=\mathrm{diag}[\mathcal{W}_{ab}(\bm{k}),-\mathcal{W}_{ab}^{*}(-\bm{k})]$ with $\mathcal{V}_{a}(\bm{k})=\partial H_{\bm{k}}/\partial k_{a}$ and $\mathcal{W}_{ab}(\bm{k})=\partial^{2}H_{\bm{k}}/\partial k_{a}\partial k_{b}$. At linear order in $\bm{q}$, we find $\delta M_{a}^{(1)}=\alpha_{ab}q_{b}$, where the first-order spin susceptibility $\alpha_{ab}$ reads
@@
where $\mathcal{G}_{0}\equiv\mathcal{G}(\bm{k},0,i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}=0})^{-1}$ for brevity. This formula is consistent with the results in Ref. Chirolli et al. (2022). Importantly, we find the nonlinear spin magnetization along $c$ direction in Eq. (1) with the second-order spin susceptibility $\chi$ as
$\chi_{ab}^{c}=-\frac{g_{s}\mu_{B}}{32\beta}\!\sum_{n\bm{k}}\!\mathrm{Tr}[\eta_{c}(\mathcal{G}_{0}\hat{w}_{ab}\mathcal{G}_{0}+2\mathcal{G}_{0}\hat{v}_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0})].$ (7)
@@
We note that that $\chi_{ab}^{c}$ in Eq. (7) is general for arbitrary model Bloch Hamiltonians as well as pairing potentials. In particular, Eq. (7) applies for both conventional pairing as well as unconventional pairing $\hat{\Delta}_{\bm{k}}$. However, in the next section we will take a simple $s$-wave pairing $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ with a $\bm{k}$-independent pairing gap as an illustrative example to emphasize the effect of the magnetic order. Detailed derivations can be found in Supplementary Material Not .
@@
It is instructive to examine the symmetry requirements for both the linear $\alpha_{ab}$ and second-order NSM $\chi_{ab}^{c}$ summarized in Table 1. For instance, $\mathcal{T}$ breaking is necessary for non-zero $\chi_{ab}^{c}$ while $\mathcal{P}$ breaking is necessary for $\alpha_{ab}$; both are destroyed in $\mathcal{PT}$-invariant systems. A particular interesting case is that of $C_{4}\mathcal{T}$ symmetry found in altermagnets: we find that $\chi_{ab}^{x}$ and $\chi_{ab}^{y}$ terms vanish; only $\chi_{ab}^{z}$ is non-zero. Interestingly, $C_{4}\mathcal{T}$ further constrains the NSM so that $\chi_{xx}^{z}=-\chi_{yy}^{z}$. In the presence of a weak SOI naturally found in realistic altermagnets, in-plane spin magnetization can be induced as expected of a linear magnetoelectric effect in Eq. (6). Importantly, however, the Rashba SOI induced linear magnetoelectric effect is purely in-plane ($\alpha_{zx}=\alpha_{zy}=0$). This means that for out-of-plane magnetatization responses, the second order NSM $\chi_{ab}^{z}$ is the leading order term in the altermagnetic systems. As we will see below, it dominates the out-of-plane magnetization responses.
@@
Figure 2: (a) The fully gapped BdG energy spectrum of ALM/SC from Eq. (2) with Bloch Hamilotnian in Eq. (8) at $k_{y}=0$ using a phenomelogical gap $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$. (b) The spin magnetization $M_{z}$ as a function of $q_{x}$ from Eq. (4) (solid orange). As a comparison, the NSM susceptibility multiplied by $q_{x}^{2}$ is plotted as Eq. (7) (dashed green). (c) The nonlinear spin susceptibility $\chi^{z}_{xx}$ as a function of $d$-wave magnetic order $J_{\mathrm{ex}}$. The inset shows the region of $0<J_{\mathrm{ex}}<0.04t$. (d) For nonzero $\lambda$, the linear and nonlinear spin susceptibilities $\alpha_{xy}$ and $\chi^{z}_{xx}$ as a function of $\lambda$. $\alpha_{xy}$ increases linearly with $\lambda$ for small $\lambda$. Parameters: $\lambda=0$ for (a), (b) and (c). $J_{\mathrm{ex}}=0.2t$ for (a), (b) and (d). $(\Delta_{0},\mu)=(0.1,0.3)t$ for all four panels. The temperature is set to be $T=0.3T_{c}$.
@@
NSM in ALM/SC heterostructure—In the set-up illustrated in Fig. 1(a), an altermagnet thin film is in contact with a conventional $s$-wave superconductor. To characterize the physics of quasi-2D planar $d_{x^{2}-y^{2}}$-wave altermagnet, we adopt a generic two-band Hamiltonian Šmejkal et al. (2022a), yielding
$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+J_{\mathrm{ex}}a^{2}(k_{x}^{2}-k_{y}^{2})s_{z}.$ (8)
@@
Here $t$ parameterizes the usual kinetic energy, $J_{\mathrm{ex}}$ denotes the $d_{x^{2}-y^{2}}$ exchange magnetic order and $\mu$ is the Fermi energy. $a$ denotes the lattice constant. We note that an additional Rashba SOI term $H_{R}=\lambda a(k_{x}s_{y}-k_{y}s_{x})$ can be readily added to $H_{\bm{k}}$ with $\lambda$ term representing the strength of SOI, that arises in altermagnets Šmejkal et al. (2022a); Fang et al. (2023). Here, $s$ matrices operate on the spin degree of freedom. This Hamiltonian breaks time reversal symmetry $\mathcal{T}=-is_{y}\mathcal{K}$ ($\mathcal{K}$ is the complex conjugate) and four-fold rotation symmetry $C_{4}=e^{i\pi s_{z}/4}$, but preserves the combination of them as $C_{4}\mathcal{T}$. The SOI term breaks inversion ($\mathcal{P}$) symmetry without breaking $C_{4}\mathcal{T}$. Throughout our work, we use the energy unit $t$. As an illustration of the spin-split band structure in altermagnet, we plot Eq. (8) in Fig. 1(c); here we have used $J_{\mathrm{ex}}=0.4t$ and $\lambda=0$.
@@
Because of the proximity effect between the altermagnet and superconductor, Cooper pairs can tunnel into the altermagnet. For a thin film of altermagnet [see Fig. 1(a)], with thickness $d$ much less than the coherence length of the superconductor ($d\ll\xi$), the pairing gap is approximately uniform along the $z$ direction. We have focused on an conventional $s$-wave superconductor with the pairing $\Delta_{0}$. Using the parameters as $(J_{\mathrm{ex}},\lambda,\mu,\Delta_{0})=(0.2,0,0.3,0.1)t$, we diagonalize the BdG Hamiltonian and plot the fully gapped BdG energy spectrum in Fig. 2(a). Although the pairing potential is purely $s$-wave, the induced superconducting correlations can be obtained as $F=[\psi_{0}+\bm{d}(\bm{k})\cdot\bm{s}]is_{y}$, where $\psi_{0}$ and $\bm{d}$ vector parameterize the spin-singlet and spin-triplet pairing correlations, respectively Gor’kov and Rashba (2001); Frigeri et al. (2004); Zhou et al. (2016). We have $d=(0,0,d_{z})$ with $d_{z}\propto J_{\mathrm{ex}}(k_{x}^{2}-k_{y}^{2})$, indicating that the $d$-wave magnetic order gives rise to the spin-triplet pairing correlation in altermagnets due to the proximity effect.

### v5

@@ Abstract
Supercurrent flow can induce a nonvanishing spin magnetization in noncentrosymmetric superconductors with spin-orbit interaction. Often known as the non-dissipative magnetoelectric effect, these are most commonly found at linear order in supercurrent flow. Here, we propose that a nonlinear superconducting magnetoelectric (NSM) effect can naturally manifest in magnet/superconductor heterostructures. In such platforms, NSM manifests as the spin polarization generated as a second-order response to a driving supercurrent. Strikingly, we find NSM survives centrosymmetry and is the leading order magnetic response in a variety of magnetic materials that include both collinear magnets [e.g., $d$-wave planar altermagnet thin film/superconductor (ALM/SC)] as well as noncollinear magnets (e.g., Kagome/SC systems). This renders NSM a powerful electric and non-dissipative means of controlling magnetization in magnet/superconductor heterostructures, a promising platform for superconducting spintronics.
Introduction.—Unlike the normal metallic state, the intrinsically non-dissipative flow of current in superconductors enables the control of spins and magnetization without dissipation Levitov et al. (1985); Edelstein (1995, 2005)—a critical functionality in realizing low-power spintronics Linder and Robinson (2015); Eschrig (2015, 2011). Microscopically, this process occurs through the superconducting magnetoelectric effect (sometimes known as the Edelstein effect) Edelstein (1995, 2005); Fujimoto (2005); Tkachov (2017); He and Law (2020, 2021); Ikeda and Yanase (2020); Chirolli et al. (2022); He et al. (2019); Yuan (2023), where a supercurrent flowing through a noncentrosymmetric metal can induce a net magnetization. In Rashba superconductors Bauer and Sigrist (2012); Smidman et al. (2017); Houzet and Meyer (2015), such magnetoelectric effects can be naturally understood through the spin-orbit interaction (SOI): spin readily couples to the motion of electrons (and their associated Cooper pairs) producing a net spin polarization in the presence of supercurrent flow.
@@
Here we unveil the nonlinear superconducting magnetoelectric (NSM) effect by examining second-order spin generation in response to the supercurrent. Specifically, we find a (second-order nonlinear) supercurrent-induced magnetization $\delta\bm{M}$ in a two dimensional platform as
@@
where $a,b=x,y$ is the direction of the supercurrent, $c=x,y,z$, and $\bm{q}$ is the Cooper pair momentum in the presence of supercurrent flow. Note $\chi_{ab}^{c}$ vanishes in the presence of $\mathcal{T}$ symmetry. As a result, to realize the second-order nonlinear response, $\mathcal{T}$ symmetry breaking is necessary. Interestingly, Eq. (1) persists in centrosymmetric systems in sharp contrast to the case of the linear magnetoelectric effect in noncentrosymmetric superconductors He and Law (2020). As we argue, NSM can be naturally realized in magnet/superconductor heterostructures by utilizing the proximity effect between a magnetic material and an $s$-wave superconductor Fu and Kane (2008); Qi et al. (2010); Wang et al. (2015); Tkachov (2017).
To be specific, we propose that NSM can dominate magnetoelectricity in a variety of magnetic material/superconductor heterostructures that include both collinear (e.g., altermagnets) as well as noncollinear (e.g., kagome lattice Lee et al. (2024a)) magnetic order. A particularly striking example is $d$-wave planar altermagnet thin film/superconductor (ALM/SC) heterostructure [see Fig. 1(a)]. Altermagnets are collinear antiferromagnets with an unconventional magnetic order Ma et al. (2021); Šmejkal et al. (2022a); Lee et al. (2024b); Mazin et al. (2022); Mazin (2023); Feng et al. (2022); Šmejkal et al. (2022b, 2023); Ghorashi et al. (2024); Fang et al. (2024); Sattigeri et al. (2023); Krempaskỳ et al. (2024); Fedchenko et al. (2024); Šmejkal et al. (2022c); Hodt and Linder (2024); Zhu et al. (2023); Beenakker and Vakhtel (2023); Zhang et al. (2024); Ouassou et al. (2023); Sun et al. (2023); Papaj (2023); Li and Liu (2023); Giil et al. (2024); Sukhachov et al. (2024); Giil and Linder (2024); Hodt et al. (2024); Zyuzin (2024) characterized by a large spin-splitting of order hundreds of millielectronvolts Šmejkal et al. (2022a). As we describe below, the large spin-splitting in altermagnets naturally produces a pronounced second-order NSM $\chi_{ab}^{z}$ that dominates the out-of-plane spin magnetization of ALM/SC heterostructures. Indeed, when its $d$-wave $C_{4}\mathcal{T}$ symmetry is preserved, we find $\chi_{ab}^{z}$ provides the leading order contribution to the out-of-plane spin magnetization sharply contrasting with polar or gyrotropic superconductors Edelstein (1995); He et al. (2019); He and Law (2020) where the linear magnetoelectric effect dominates. Candidate altermagnetic materials include RuO2 and KRu4O8 thin films that can produce sizeable NSM susceptibilities.
@@
General theory of nonlinear superconducting magnetoelectric effect.—We first examine the superconducting magnetoelectric effect within an effective Bogoliubov-de Gennes (BdG) framework. We note that our treatment is general and is agnostic to the precise microscopics of the superconducting states, which can be intrinsic or proximity induced. In what follows, we describe either case phenomenologically via the pairing potential $\hat{\Delta}_{\bm{k}}$. The finite-$\bm{q}$ BdG Hamiltonian in the Nambu basis $\Psi_{\bm{k},\bm{q}}=(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})^{T}$ reads
@@
where $H_{\bm{k}}$ is the Bloch Hamiltonian in the normal state, and $\bm{q}$ is the momentum of the Cooper pair. For simplicity we have used a $\bm{q}$-independent gap function $\hat{\Delta}_{\bm{k}}$, which is valid for weak values of the applied current Chirolli et al. (2022).
The free energy $\mathcal{F}$ of the heterostructure [Fig. 1(a)] comprises two parts: the superconductor component $\mathcal{F}_{\mathrm{sc}}$ and the material (i.e. altermagnet) component $\mathcal{F}_{m}$ so that $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{m}$. The supercurrent can be tracked as $\bm{J}_{s}=\nu_{s}\bm{q}$ with the condensate superfluid stiffness $\nu_{s}=\partial^{2}\mathcal{F}_{\mathrm{sc}}(\bm{q})/\partial^{2}\bm{q}$. As a result, observables that scale with $q$ below are induced by the supercurrent at linear order while $q^{2}$ indicates a second-order nonlinear response.
To obtain the magnetic response, we analyze the set of parametric partition function $\mathcal{Z}(\bm{h})$
@@
where $\bm{h}$ is a parameter describing an auxiliary Zeeman field with a corresponding Zeeman energy $\Sigma(\bm{h})=\mathrm{diag}(g_{s}\mu_{B}\bm{h}\cdot\hat{s},-g_{s}\mu_{B}\bm{h}\cdot\hat{s}^{*})/2$. Here $g_{s}=2$ is the Landé $g$ factor, $\mu_{B}$ is the Bohr magneton, and the Gor’kov’s Green’s functions are $G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})=\mathcal{G}^{-1}(\bm{k},\bm{q},i\omega_{n})+\Sigma(\bm{h})$ and $\mathcal{G}(\bm{k},\bm{q},i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}})^{-1}$ with the Matsubara frequency $\omega_{n}=(2n+1)\pi/\beta$. $\beta=1/k_{B}T$ with $T$ being the temperature.
@@
Here $\eta_{a}=\mathrm{diag}(s_{a},-s_{a}^{*})$ are the generalized spin Pauli matrices in the Nambu space. We then extract the linear and nonlinear responses systematically, and at linear order in $\bm{q}$ we find $\delta M_{a}^{(1)}=\alpha_{ab}q_{b}$, where the first-order spin susceptibility $\alpha_{ab}$ reads
@@
where $\mathcal{G}_{0}\equiv\mathcal{G}(\bm{k},0,i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}=0})^{-1}$ for brevity. Here $\hat{v}_{a}=\mathrm{diag}[\mathcal{V}_{a}(\bm{k}),-\mathcal{V}_{a}^{*}(-\bm{k})]$ with $\mathcal{V}_{a}(\bm{k})=\partial H_{\bm{k}}/\partial k_{a}$. This formula is consistent with the results in Ref. Chirolli et al. (2022). Importantly, we find the nonlinear spin magnetization along $c$ direction in Eq. (1) with the second-order spin susceptibility $\chi$ as
$\chi_{ab}^{c}=-\frac{g_{s}\mu_{B}}{32\beta}\!\sum_{n\bm{k}}\!\mathrm{Tr}[\eta_{c}(\mathcal{G}_{0}\hat{w}_{ab}\mathcal{G}_{0}+2\mathcal{G}_{0}\hat{v}_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0})],$ (6)
@@
where $\hat{w}_{ab}=\mathrm{diag}[\mathcal{W}_{ab}(\bm{k}),-\mathcal{W}_{ab}^{*}(-\bm{k})]$ with $\mathcal{W}_{ab}(\bm{k})=\partial^{2}H_{\bm{k}}/\partial k_{a}\partial k_{b}$. We note that $\chi_{ab}^{c}$ in Eq. (6) is general for various model Bloch Hamiltonians and applies for both conventional pairing as well as unconventional pairing $\hat{\Delta}_{\bm{k}}$. In the following we will take a simple $s$-wave pairing $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ with a $\bm{k}$-independent pairing gap as an illustrative example to emphasize the effect of the magnetic order. More details of the derivations can be found in Supplementary Material (SM) Not .
Table 1: Symmetry restrictions for three components of $\bm{M}$ in a two-dimensional system. Note that $C_{4}\mathcal{T}$ corresponds to the case for altermagnets that we focus on in our work. Centrosymmetric $D_{3d}$ group corresponds to the kagome lattice. 1(2) indicates first (second) order response. $\checkmark$ indicates allowed and $\times$ indicates forbidden.
$\mathcal{P}$ $\mathcal{T}$ $\mathcal{PT}$ $C_{4}\mathcal{T}$ $D_{3d}$
$M_{x}$(2) ✓(1) $\times$ ✓(1) ✓(2)
$M_{y}$(2) ✓(1) $\times$ ✓(1) ✓(2)
$M_{z}$(2) ✓(1) $\times$ ✓(2) $\times$
@@
It is instructive to examine the symmetry requirements for both the linear $\alpha_{ab}$ and second-order NSM $\chi_{ab}^{c}$ summarized in Table 1. For instance, $\mathcal{T}$ breaking is necessary for non-zero $\chi_{ab}^{c}$ while $\mathcal{P}$ breaking is necessary for $\alpha_{ab}$; both are destroyed in $\mathcal{PT}$-invariant systems. A particular interesting case is that of $C_{4}\mathcal{T}$ symmetry found in altermagnets: we find that $\chi_{ab}^{x}$ and $\chi_{ab}^{y}$ terms vanish; only $\chi_{ab}^{z}$ is non-zero. Interestingly, $C_{4}\mathcal{T}$ further constrains the NSM so that $\chi_{xx}^{z}=-\chi_{yy}^{z}$. In the presence of a weak SOI naturally found in realistic altermagnets, in-plane spin magnetization can be induced as expected of a linear magnetoelectric effect in Eq. (5). Importantly, the Rashba SOI induced linear magnetoelectric effect is purely in-plane ($\alpha_{zx}=\alpha_{zy}=0$). This means that for out-of-plane magnetic responses, the second order NSM $\chi_{ab}^{z}$ is the leading order term in the altermagnetic systems. As we will see below, it dominates the out-of-plane magnetic responses. We note that Ref. Zyuzin (2024) recently discussed how $d$-wave magnetic order in a superconductor induces a second-order NSM response. Our general formulation reveals that such NSM responses also extend beyond altermagnets. An illustrative example involves materials in the $D_{3d}$ group; in such systems, we find that NSM persists and dominates the in-plane magnetic responses, see below.
@@
Figure 2: (a) The BdG energy spectrum of ALM/SC from Eq. (2) with Bloch Hamilotnian in Eq. (7) at $k_{y}=0$ with spin-singlet pairing $\Delta_{0}$. (b) The spin magnetization $M_{z}$ as a function of $q_{x}$ from Eq. (4) (solid orange). As a comparison, the NSM susceptibility multiplied by $q_{x}^{2}$ is plotted as Eq. (6) (dashed green). (c) The nonlinear spin susceptibility $\chi^{z}_{xx}$ as a function of $d$-wave magnetic order $J_{\mathrm{ex}}$. The inset shows the region of $0<J_{\mathrm{ex}}<0.04t$. (d) For nonzero $\lambda$, the linear and nonlinear spin susceptibilities $\alpha_{xy}$ and $\chi^{z}_{xx}$ as a function of $\lambda$. $\alpha_{xy}$ increases linearly with $\lambda$ for small $\lambda$. Parameters: $\lambda=0$ for (a), (b) and (c). $J_{\mathrm{ex}}=0.2t$ for (a), (b) and (d). $(\Delta_{0},\mu)=(0.1,0.3)t$ for all four panels. The temperature is set to be $T=0.3T_{c}$. In our calculations, we adopt the BCS temperature dependence of $\Delta_{0}$ with $\Delta(T)=\Delta_{0}\tanh(1.74\sqrt{T_{c}/T-1})$ and $\Delta_{0}=1.76k_{B}T_{c}$.
@@
NSM in ALM/SC heterostructure.—To exemplify NSM, we first focus on the ALM/SC set-up illustrated in Fig. 1(a): an altermagnet thin film is in contact with a conventional $s$-wave superconductor. To characterize the physics of quasi-2D planar $d_{x^{2}-y^{2}}$-wave altermagnet, we adopt a generic two-band Hamiltonian Šmejkal et al. (2022a), yielding
$H_{\bm{k}}=(ta^{2}\bm{k}^{2}-\mu)s_{0}+J_{\mathrm{ex}}a^{2}(k_{x}^{2}-k_{y}^{2})s_{z}.$ (7)
@@
Here $t$ parameterizes the usual kinetic energy, $J_{\mathrm{ex}}$ denotes the $d_{x^{2}-y^{2}}$ exchange magnetic order and $\mu$ is the Fermi energy. $a$ denotes the lattice constant. We note that an additional Rashba SOI term $H_{R}=\lambda a(k_{x}s_{y}-k_{y}s_{x})$ can be readily added to $H_{\bm{k}}$ with $\lambda$ term representing the strength of SOI, that arises in altermagnets Šmejkal et al. (2022a); Fang et al. (2024). This Hamiltonian breaks time reversal symmetry $\mathcal{T}=-is_{y}\mathcal{K}$ ($\mathcal{K}$ is the complex conjugate) and four-fold rotation symmetry $C_{4}=e^{i\pi s_{z}/4}$, but preserves the combination of them as $C_{4}\mathcal{T}$. The SOI term breaks inversion ($\mathcal{P}$) symmetry without breaking $C_{4}\mathcal{T}$. Throughout our work, we use the energy unit $t$. As an illustration of the spin-split band structure in altermagnet, we plot Eq. (7) in Fig. 1(c); here we have used $J_{\mathrm{ex}}=0.4t$ and $\lambda=0$.
@@
Because of the proximity effect between the altermagnet and superconductor, Cooper pairs can tunnel into the altermagnet. For a thin film of altermagnet [see Fig. 1(a)], with thickness $d$ much less than the coherence length of the superconductor ($d\ll\xi$), the pairing gap is approximately uniform along the $z$ direction. We have focused on an conventional $s$-wave superconductor with the pairing $\Delta_{0}$. Using the parameters as $(J_{\mathrm{ex}},\lambda,\mu,\Delta_{0})=(0.2,0,0.3,0.1)t$, we diagonalize the BdG Hamiltonian and plot the fully gapped BdG energy spectrum in Fig. 2(a). Although the pairing potential is purely $s$-wave, the induced superconducting correlations can be obtained as $F(\bm{k},\omega)=[\psi_{0}+i\omega\bm{d}(\bm{k})\cdot\bm{s}]is_{y}$, where $\psi_{0}$ and $\bm{d}$ vector parameterize the spin-singlet and spin-triplet pairing correlations, respectively Bergeret et al. (2005); Gor’kov and Rashba (2001); Frigeri et al. (2004); Zhou et al. (2016); Linder and Balatsky (2019). We have $d=(0,0,d_{z})$ with $d_{z}\propto J_{\mathrm{ex}}(k_{x}^{2}-k_{y}^{2})$, indicating that the $d$-wave magnetic order gives rise to the odd-frequency spin-triplet pairing correlation in altermagnets due to the proximity effect.